OpenStudy (anonymous):

Find the real-number root. sqrt -2.56

OpenStudy (jdoe0001):

\(\bf \sqrt{-2.56}\implies \sqrt{-1\cdot 2.56}\implies \sqrt{-1}\cdot \sqrt{2.56}\implies \sqrt{2.56}\ i\implies ?\)

OpenStudy (anonymous):

1.13137085 + 1.13137085 i

OpenStudy (anonymous):

It has no real root number?

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

well I'm thinking you're asked for the positive factor of it

OpenStudy (jdoe0001):

\(\bf \sqrt{-2.56}\implies \sqrt{-1\cdot 2.56}\implies \sqrt{-1}\cdot \sqrt{2.56}\implies {\color{brown}{ \sqrt{2.56}}}\ i\implies {\color{brown}{ 1.6}}\ i\)

OpenStudy (jdoe0001):

but yes, the radicand is negative, and the root is even, thus is an imaginary value, so no real root for the whole

OpenStudy (anonymous):

Okay thanks :)